# Stoichiometry Representation Learning with Polymorphic Crystal Structures

### v1

@@ 6 Conclusion
This paper focuses on learning a probabilistic representation of stoichiometry that incorporates polymorphic structural information of crystalline materials. Given stoichiometry and its corresponding polymorphic structures, PolySRL learns parameterized Gaussian distribution for each stoichiometry, whose mean becomes the representation of stoichiometry and variance indicates the level of uncertainty stemming from the polymorphic structures. Extensive empirical studies on sixteen datasets, including wet-lab experimental data and DFT-calculated data, have been conducted to validate the effectiveness of PolySRL in learning stoichiometry representations. Moreover, a comprehensive analysis of uncertainties reveals that the model learns diverse complexities encountered in materials science, highlighting the practicality of PolySRL in real-world material discovery process.
@@
Stoichiometry Representation Learning with Polymorphic Crystal Structures
@@
This is an Appendix for the paper Stoichiometry Representation Learning with Polymorphic Crystal Structures, which is organized as follows: Section A provides details on constructing structural graph representation of crystalline materials. Section B elaborates on the implementation details of our method. Section C details all the datasets we use. Section D details the experimental setup of all the baseline methods. Section E describes evaluation protocol. Section F provides additional experimental results. Section G lists important notations used during the main manuscript.
@@ Appendix A Structural Graph Construction
In this section, we provide the detailed structural graph construction process with a figure. Overall, this structural graph is the same as previous works Xie & Grossman (2018); Yan et al. (2022). Given a crystal structure $(\mathbf{P},\mathbf{L})$, suppose the unit cell has $n_{s}$ atoms, we have $\mathbf{P}={[\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{n_{s}}]}^{\intercal}\in\mathbb{R}^{{n_{s}}\times 3}$ indicating the atom position matrix and $\mathbf{L}={[\mathbf{l}_{1},\mathbf{l}_{2},\mathbf{l}_{3}]}^{\intercal}\in\mathbb{R}^{3\times 3}$ representing the lattice parameter describing how a unit cell repeats itself in three directions. Since the crystal usually possesses irregular shapes in practice, $\mathbf{l}_{1},\mathbf{l}_{2},\mathbf{l}_{3}$ are not always orthogonal in 3D space Yan et al. (2022). For clear visualization, we provide examples of periodic patterns in 2D space in Figure 5 (a).
Based on the crystal parameters $(\mathbf{P},\mathbf{L})$, we construct a multi-edge graph $\mathcal{G}^{b}=(\mathcal{V},\mathbf{A}^{b})$ that captures atom interactions across cell boundaries Xie & Grossman (2018). Specifically, $v_{i}\in\mathcal{V}$ denotes an atom $i$ and all its duplicates in the infinite 3D space whose positions are included in the set $\{\hat{\mathbf{p}}_{i}|\hat{\mathbf{p}}_{i}=\mathbf{p}_{i}+k_{1}\mathbf{l}_{1}+k_{2}\mathbf{l}_{2}+k_{3}\mathbf{l}_{3},k_{1},k_{2},k_{3}\in\mathbb{Z}\}$, where $\mathbb{Z}$ denotes the set of all the integers. Moreover, $\mathbf{A}^{b}\in\{0,1\}^{n_{s}\times n_{s}}$ denotes an adjacency matrix, where $\mathbf{A}_{i,j}^{b}=1$ if two atoms $i$ and $j$ are within the predefined radius $r$ and $\mathbf{A}^{b}_{ij}=0$ otherwise. Specifically, nodes $v_{i}$ and $v_{j}$ are connected if there exists any combination $k_{1},k_{2},k_{3}\in\mathbb{Z}$ such that the euclidean distance $d_{ij}$ satisfies $d_{ij}={\|\mathbf{p}_{i}+k_{1}\mathbf{l}_{1}+k_{2}\mathbf{l}_{2}+k_{3}\mathbf{l}_{3}-\mathbf{p}_{j}\|}_{2}\leq r$ (see Figure 5 (b)). For the initial feature for edges, we expand the distance $d_{ij}$ between atom $v_{i}$ and $v_{j}$ by Gaussian basis following previous works Xie & Grossman (2018). Moreover, for each node in $\mathcal{G}^{b}$ is associated with a learnable feature $\mathbf{x}^{b}\in\mathbb{R}^{F}$, which is shared across all crystals, to make sure we utilize only structural information.
@@ Appendix B Implementation Details
In this section, we provide implementation details of PolySRL.
@@ B.1 Structural Graph Encoder
$\mathbf{h}^{0,b}_{i}=\phi_{node}(\mathbf{X}^{b}),~{}~{}~{}\mathbf{b}^{0,b}_{ij}=\phi_{edge}(\mathbf{B}^{b}_{ij}),$ (7)
where $\mathbf{X}^{b}\in\mathbb{R}^{n_{s}\times F}$ is the atom feature matrix whose $i$-th row indicates the input feature of atom $v_{i}$, $\mathbf{B}^{b}\in\mathbb{R}^{n_{s}\times n_{s}\times F}$ is the bond feature tensor. As previously explained in Section 3.2, we employ a common $\mathbf{x}^{b}$ for all atoms across all crystals, resulting in every row in $\mathbf{X}^{b}$ being identical to $\mathbf{x}^{b}$. With the initial representations of atoms and bonds, the processor learns to pass messages across the crystal structure and update atom and bond representations as follows:
$\mathbf{b}^{l+1,b}_{ij}=\psi^{l}_{edge}(\mathbf{h}^{l,b}_{i},\mathbf{h}^{l,b}_{j},\mathbf{b}^{l,b}_{ij}),~{}~{}~{}\mathbf{h}^{l+1,b}_{i}=\psi^{l}_{node}(\mathbf{h}^{l,b}_{i},\sum_{j\in\mathcal{N}(i)}{\mathbf{b}^{l+1,b}_{ij}}),$ (8)
@@
B.2 Probabilistic Stoichiometry Encoder
Stoichiometry Graph Encoder $f^{a}$. For the stoichiometry graph encoder $f^{a}$, we utilize the architecture of GCNs Kipf & Welling (2016) and Jumping Knowledge Network Xu et al. (2018). Specifically, given elemental feature matrix $\mathbf{X}^{a}$ and adjacency $\mathbf{A}^{a}$, GCN layers pass the messages to obtain latent elemental feature matrix as follows:
@@ B.2 Probabilistic Stoichiometry Encoder
where $\mathbf{h}^{0,a}_{i}$ indicates $i$-th row of elemental feature matrix $\mathbf{X}^{a}$, and $l=0,\ldots,L^{\prime}$. After $L^{\prime}$ step message passing steps, we obtain a final representation of stoichiometry as follows:
@@
where $\mathbf{W}\in\mathbb{R}^{F\times L^{\prime}F}$ is a learnable weight matrix that reduces the dimension of concatenated representations. Note that $\mathbf{Z}^{a}_{i}$ is equivalent to the $i$-th row of the element embedding matrix $\mathbf{Z}^{a}$ in Equation 2. We also use $L^{\prime}=3$ for stoichiometry encoder $f^{a}$. After obtaining the elemental representation matrix $\mathbf{Z}^{a}$, we obtain stoichiometry representation $\mathbf{z}^{a}$ by employing weighted sum pooling, which takes into account the compositional ratio.
Mean $f_{\mu}^{a}$ and Variance $f_{\sigma}^{a}$ Module. After obtaining the elemental representation matrix $\mathbf{Z}^{a}$, we utilize set2set Vinyals et al. (2015) pooling function to obtain the mean vector and diagonal entries of the covariance vector. More specifically, given $\mathbf{Z}^{a}$, we obtain mean vector $\mathbf{z}^{a}_{\mu}$ and diagonal covariance vector $\mathbf{z}^{a}_{\sigma}$ as follows:
$\mathbf{z}^{a}_{\mu}=\mathbf{\hat{z}}^{a}_{\mu}+\mathbf{z}^{a},~{}~{}~{}\mathbf{\hat{z}}^{a}_{\mu}=\text{Set2set}_{\mu}(\mathbf{Z}^{a}),$ (11)
$\mathbf{z}^{a}_{\sigma}=\mathbf{\hat{z}}^{a}_{\sigma}+\mathbf{z}^{a},~{}~{}~{}\mathbf{\hat{z}}^{a}_{\sigma}=\text{Set2set}_{\sigma}(\mathbf{Z}^{a}).$ (12)
@@ B.3 Training Details
We also describe the implementation details to enhance the reproducibility. Our method is implemented on Python 3.7.1, PyTorch 1.8.1, and Torch-geometric 1.7.0. All experiments are conducted using a 24GB NVIDIA GeForce RTX 3090. Model hyperparameters are given in Table 2. During training, we clip the gradient to the maximum value of 2 for stability Zhang et al. (2019).
Table 2: Hyperparameter specifications of PolySRL.
@@
Appendix C Datasets
  • Materials Project Jain et al. (2013) is an openly accessible database that provides material properties calculated using density functional theory (DFT). We have gathered 80,162 distinct stoichiometries along with their corresponding 112,183 crystal structures computed using DFT, with up to 32,021 stoichiometries having multiple potential structures.
  • Band Gap Zhuo et al. (2018) dataset comprises experimentally determined band gap properties for non-metallic materials. It encompasses 2,482 distinct stoichiometries and a total of 3,895 experimental band gap values. Within this dataset, 1,413 instances of duplicate experimental band gap measurements for stoichiometries were identified. Consequently, our task involves predicting the band gap properties for these 2,482 stoichiometries, with the average value being computed in cases where duplicate experimental results exist for a given stoichiometry.
  • Formation Enthalpies Kim et al. (2017) dataset consists of experimentally determined formation enthalpy values for intermetallic phases and other inorganic compounds. It includes 1,141 unique stoichiometries and a total of 1,276 experimental formation enthalpy values. Within this dataset, 135 cases of duplicate experimental formation enthalpy measurements for stoichiometries were identified. Therefore, our objective is to predict the formation enthalpy properties for these 1,141 stoichiometries, calculating the average value when duplicate experimental results are present for a particular stoichiometry. We report MAE values multiplied by a factor of 10 for clear interpretation during all experiments.
  • Metallic Morgan (2018) dataset contains reduced glass transition temperature (Trg) for 584 unique metallic alloys. We report MAE values multiplied by a factor of 10 for clear interpretation during all experiments.
  • ESTM 300 K Na & Chang (2022) dataset contains various properties of 368 thermoelectric materials that are measured in the temperature range of 295 K to 305 K, which is widely recognized as room temperature in chemistry. Among the properties, we mainly target electrical conductivity ($S/m$), thermal conductivity ($W/mK$), and Seebeck coefficient ($\mu V/K$). Regarding electrical conductivity and thermal conductivity, we apply a logarithmic scaling to the target values because they exhibit significant skewness. Additionally, for the Seebeck coefficient, we use min-max scaling on the target values due to their wide range and report MAE values multiplied by a factor of 10 for clear interpretation during all experiments. When calculating the figure of merit ($Z\bar{T}$) with predicted properties, we reverse the scaling to return the original scale and then compute it.
  • ESTM 600 K Na & Chang (2022) dataset contains various properties of 188 thermoelectric materials that are measured in the temperature range of 593 K to 608 K, which is widely recognized as high temperature in chemistry. The properties we are targeting and the preprocessing steps applied are identical to those used for the ESTM 300 K dataset.
In addition to the wet-lab experimental datasets, we use following seven Matbench datasets that contain properties from DFT calculation.
  • [leftmargin=.1in]
  • Castelli Perovskites Castelli et al. (2012) dataset contains formation energy of Perovskite cell of 18,928 materials.
  • Refractive Index Jain et al. (2013) dataset contains a refractive index of 4,764 materials, provided in MP database.
  • Shear Modulus Jain et al. (2013) dataset contains shear modulus of 10,987 materials, provided in MP database.
  • Bulk Modulus Jain et al. (2013) dataset contains bulk modulus of 10,987 materials, provided in MP database.
  • Exfoliation Energy Choudhary et al. (2017) dataset contains exfoliation energy 636 materials.
  • MP Band gap Jain et al. (2013) dataset contains band gap of 106,113 materials, provided in MP database.
  • MP Formation Energy Jain et al. (2013) dataset contains formation energy per atom in 132,752 materials, provided in MP database.
Following previous work Wang et al. (2021); Goodall & Lee (2020), we choose the target value associated with the lowest formation enthalpy for duplicate stoichiometries found in both the MP datasets, while we use the mean of the target values for other datasets.
Appendix D Baseline Methods
  • Rand init. refers to the randomly initialized stoichiometry encoder without any training process.
  • GraphCL You et al. (2020) is a general graph-level contrastive learning strategy that uses random augmentation to construct positive and negative samples. In this paper, it learns the stoichiometry representation based on the random augmentation on the stoichiometry graph $\mathcal{G}^{a}$, without utilizing structural information. For the $n$-th data in the minibatch ($N$ data points), the loss function is defined as follows follows:
  • MP Band G. and MP Form. E. learn the stoichiometry representation by predicting the DFT-calculated properties, i.e., band gap and formation energy per atom, respectively. More formally, model is trained with MAE loss for $n$-th data point in the minibatch ($N$ data points) as follows:
    $l_{n}=|Y_{n}-\hat{Y}_{n}|,\$ (14)
  • 3D Infomax Stärk et al. (2022) proposes to enhance model prediction on 3D molecular graphs by integrating 3D information of the molecules in its latent representations. Instead of 2D molecular graphs, we learn the representation of stoichiometry graph $\mathcal{G}^{a}$ by maximizing the mutual information with structural graph $\mathcal{G}^{b}$. More specifically, we train the model with NTXent (Normalized Temperature-scaled Cross Entropy) loss Chen et al. (2020), which is defined for $n$-th data point in minibatch of size $N$ as follows:
Even though the primary focus of this paper is to introduce training strategies for stoichiometry encoders without any label information, we also conduct a comparative analysis of our proposed approach with previous supervised stoichiometry representation learning methods Goodall & Lee (2020); Wang et al. (2021). Note that these works propose sophisticated model architectures for stoichiometry representation learning, not training strategy.
  • Roost Goodall & Lee (2020) first proposes to utilize GNNs for stoichiometry representation learning by presenting stoichiometry as a fully connected graph, whose nodes are unique elements in stoichiometry. This approach allows the model to acquire distinct and material-specific representations for each element, enabling it to capture physically meaningful properties and interactions.
  • CrabNet Wang et al. (2021) designs a Transformer self-attention mechanism Vaswani et al. (2017) to adaptively learn the representation of individual elements based on their chemical environment.
Appendix E Evaluation Protocol
Evaluation Metrics. We mainly compare the methods in terms of Mean Absolute Error (MAE) following previous work Goodall & Lee (2020). Moreover, we provide the model performance in terms of $R^{2}$ in Appendix F, which provides an intuitive measure of the fraction of the overall variance in the data that the model can account for.
Representation Learning. For representation learning scenarios, we fix the model parameters (i.e., $f^{a}$, $f_{\mu}^{a}$, and $f_{\sigma}^{a}$) and train a three-layer MLP head with LeakyReLU non-linearity to evaluate the stoichiometry obtained by various models. Following previous works Veličković et al. (2018); You et al. (2020), we train the MLP head with Adam optimizer with a fixed learning rate of 0.001 for 300 epochs.
Appendix F Additional Experiments
F.1 Experiments on DFT-Calculated Datasets
Although DFT-calculated properties frequently differ from actual wet-lab experimental properties Jha et al. (2019), we have included experimental outcomes for seven DFT-calculated properties from the Matbench dataset Dunn et al. (2020). These Matbench datasets were assessed using a five-fold cross-validation approach with train/validation/test splits set at a ratio of 72/8/20, as given in previous work Wang et al. (2021). In Table 3, we have following observations: 1) In the DFT-based dataset, we observed significant disparities in trends compared to the experimental datasets in Table 1, demonstrating the inherent difference between the experimental data and DFT-calculated data. For instance, we noticed that the MP Form. E. model consistently outperforms the MP Band G. and 3D Infomax models. 2) Furthermore, given that the datasets are designed to pick the target value linked to the lowest formation enthalpy among different polymorphic structures for a single stoichiometry, we find that models trained with specific DFT-calculated values (i.e., Prop. ) do not outperform models trained on corresponding datasets. This discrepancy is attributed to properties derived from non-lowest formation enthalpy polymorphic structures, which can introduce confusion to the model. 3) However, we observe PolySRL generally outperforms baseline models, demonstrating its effectiveness in not only wet-lab experimental datasets but also in DFT-calculated datasets.
Model DFT Poly. Castelli Refractive Shear Bulk Exfoliation MP
Prop. Str. Perovskites Index Modulus Modulus Energy Band G. Form. E.
Rand init. 0.140 0.394 0.115 0.850 0.393 0.354 0.119
(0.004) (0.091) (0.003) (0.030) (0.044) (0.005) (0.002)
GraphCL 0.145 0.386 0.117 0.844 0.411 0.351 0.121
(0.006) (0.094) (0.002) (0.021) (0.060) (0.004) (0.001)
MP Band G. 0.141 0.399 0.116 0.851 0.397 0.354 0.119
(0.004) (0.085) (0.002) (0.042) (0.041) (0.007) (0.002)
MP Form. E. 0.134 0.379 0.108 0.801 0.382 0.338 0.115
(0.004) (0.093) (0.002) (0.029) (0.037) (0.002) (0.001)
3D Infomax 0.147 0.388 0.117 0.880 0.408 0.354 0.116
(0.004) (0.094) (0.003) (0.040) (0.043) (0.005) (0.002)
PolySRL 0.132 0.394 0.107 0.837 0.378 0.328 0.112
(0.007) (0.092) (0.002) (0.033) (0.021) (0.005) (0.003)
F.2 Transfer Learning
In addition to evaluations on the learned stoichiometry representations, we also compare the models’ performance in transfer learning scenarios in Table 4, where the encoder parameters are fine-tuned along with the MLP head. We have the following observations: 1) Although the overall performance enhancement is observed due to the additional training of the encoder when compared with the results reported in Table 1, we sometimes observe that negative transfer occurs when comparing the Rand init. model and baseline methods in Table 4. This indicates that without an elaborate design of the tasks, pre-training may incur negative knowledge transfer to the downstream tasks Zhang et al. (2022). 2) However, by comparing to Rand init. in Table 4, we observe that PolySRL consistently leads to positive transfer to the model. We attribute this to the probabilistic representation, which maintains a high variance for uncertain materials, thereby preventing the representations of the materials from overfitting to the pretraining task.
Given that the primary objective of this paper is to propose a training approach for stoichiometry representation learning rather than introducing a new model architecture, previous supervised learning methods, i.e., Roost Goodall & Lee (2020) and CrabNet Wang et al. (2021), are not directly relevant to our research. Nevertheless, we include a comparison with these previous works in this section to offer additional insights into our model’s performance. For the experiment, we used publicly available codes provided by the authors  [^5] Roost: https://zenodo.org/record/4133793  [^6] CrabNet: https://github.com/anthony-wang/CrabNet. In Table 4, we observe that our simple stoichiometry encoder composed of GCNs and Jumping Knowledge Network (i.e., Rand init.) exhibits comparable or superior performance compared to the previous works that are elaborately designed for supervised stoichiometry learning. While previous works are elaborately designed for predicting properties of stoichiometry, they train the models from large-scale DFT-calculated properties of lowest enthalpy structures, giving up polymorphic structures of a single stoichiometry. However, in real-world scenarios, large-scale wet-lab experimental data is seldom available, which restricts their utility in the materials discovery process.
Table 4: Transfer learning performance including supervised learning baselines (MAE).

### v2

@@ 4.6 Conclusion
This paper focuses on learning a probabilistic representation of composition that incorporates polymorphism of crystalline materials. Given composition and its corresponding polymorphic structures, PCRL learns a parameterized Gaussian distribution for each composition, whose mean becomes the representation of composition and variance indicates the level of uncertainty stemming from the polymorphic structures. Extensive empirical studies on sixteen datasets have been conducted to validate the effectiveness of PCRL in learning compositional representations. Moreover, a comprehensive analysis of uncertainties reveals that the model learns diverse complexities encountered in materials science, highlighting the practicality of PCRL in real-world material discovery process.
Intended Users
PCRL is intended for chemistry, material science, and artificial intelligence researchers and data scientists who want to apply AI algorithms and innovate novel methods to tackle problems formulated in PCRL datasets and tasks.
Code Availability
The source code for PCRL is available at https://github.com/Namkyeong/PCRL.
Computing Resources
We perform all pre-training and downstream tasks on a 24GB NVIDIA GeForce RTX 3090 GPU.
Ethics Statement
The development and dissemination of the PCRL adhere to stringent ethical standards to ensure the responsible use of the data. The sources of the data are clearly documented in Section 2.5. Since this paper focuses on material discovery, we believe there are no additional ethical concerns to declare.
Competing Interests
The authors declare no competing interests.
@@
Compositional Representation of Polymorphic Crystalline Materials
@@ Appendix A Structural Graph Construction
In this section, we provide the detailed structural graph construction process with a figure. Overall, this structural graph is the same as previous works [3, 71]. Given a crystal structure $(\mathbf{P},\mathbf{L})$, suppose the unit cell has $n_{s}$ atoms, we have $\mathbf{P}={[\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{n_{s}}]}^{\intercal}\in\mathbb{R}^{{n_{s}}\times 3}$ indicating the atom position matrix and $\mathbf{L}={[\mathbf{l}_{1},\mathbf{l}_{2},\mathbf{l}_{3}]}^{\intercal}\in\mathbb{R}^{3\times 3}$ representing the lattice parameter describing how a unit cell repeats itself in three directions. Since the crystal usually possesses irregular shapes in practice, $\mathbf{l}_{1},\mathbf{l}_{2},\mathbf{l}_{3}$ are not always orthogonal in 3D space [71]. For clear visualization, we provide examples of periodic patterns in 2D space in Figure 6 (a).
Based on the crystal parameters $(\mathbf{P},\mathbf{L})$, we construct a multi-edge graph $\mathcal{G}^{b}=(\mathcal{V},\mathbf{A}^{b})$ that captures atom interactions across cell boundaries [3]. Specifically, $v_{i}\in\mathcal{V}$ denotes an atom $i$ and all its duplicates in the infinite 3D space whose positions are included in the set $\{\hat{\mathbf{p}}_{i}|\hat{\mathbf{p}}_{i}=\mathbf{p}_{i}+k_{1}\mathbf{l}_{1}+k_{2}\mathbf{l}_{2}+k_{3}\mathbf{l}_{3},k_{1},k_{2},k_{3}\in\mathbb{Z}\}$, where $\mathbb{Z}$ denotes the set of all the integers. Moreover, $\mathbf{A}^{b}\in\{0,1\}^{n_{s}\times n_{s}}$ denotes an adjacency matrix, where $\mathbf{A}_{i,j}^{b}=1$ if two atoms $i$ and $j$ are within the predefined radius $r$ and $\mathbf{A}^{b}_{ij}=0$ otherwise. Specifically, nodes $v_{i}$ and $v_{j}$ are connected if there exists any combination $k_{1},k_{2},k_{3}\in\mathbb{Z}$ such that the euclidean distance $d_{ij}$ satisfies $d_{ij}={\|\mathbf{p}_{i}+k_{1}\mathbf{l}_{1}+k_{2}\mathbf{l}_{2}+k_{3}\mathbf{l}_{3}-\mathbf{p}_{j}\|}_{2}\leq r$ (see Figure 6 (b)). For the initial feature for edges, we expand the distance $d_{ij}$ between atom $v_{i}$ and $v_{j}$ by Gaussian basis following previous works [3]. Moreover, each node in $\mathcal{G}^{b}$ is associated with a learnable feature $\mathbf{x}^{b}\in\mathbb{R}^{F}$, which is shared across all crystals, to make sure we utilize only structural information.
@@ Appendix B Implementation Details
In this section, we provide implementation details of PCRL.
@@ B.1 Structural Graph Encoder
$\mathbf{h}^{0,b}_{i}=\phi_{node}(\mathbf{X}^{b}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{b}^{0,b}_{ij}=\phi_{edge}(\mathbf{B}^{b}_{ij}),$ (10)
where $\mathbf{X}^{b}\in\mathbb{R}^{n_{s}\times F}$ is the atom feature matrix whose $i$-th row indicates the input feature of atom $v_{i}$, $\mathbf{B}^{b}\in\mathbb{R}^{n_{s}\times n_{s}\times F}$ is the bond feature tensor. As previously explained in Section 2.1.2, we employ a common $\mathbf{x}^{b}$ for all atoms across all crystals, resulting in every row in $\mathbf{X}^{b}$ being identical to $\mathbf{x}^{b}$. With the initial representations of atoms and bonds, the processor learns to pass messages across the crystal structure and update atom and bond representations as follows:
$\mathbf{b}^{l+1,b}_{ij}=\psi^{l}_{edge}(\mathbf{h}^{l,b}_{i},\mathbf{h}^{l,b}_{j},\mathbf{b}^{l,b}_{ij}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{h}^{l+1,b}_{i}=\psi^{l}_{node}(\mathbf{h}^{l,b}_{i},\sum_{j\in\mathcal{N}(i)}{\mathbf{b}^{l+1,b}_{ij}}),$ (11)
@@
B.2 Probabilistic Composition Encoder
Composition Graph Encoder $f^{a}$. For the composition graph encoder $f^{a}$, we utilize the architecture of GCNs [26] and Jumping Knowledge Network [27]. Specifically, given elemental feature matrix $\mathbf{X}^{a}$ and adjacency $\mathbf{A}^{a}$, GCN layers pass the messages to obtain latent elemental feature matrix as follows:
@@ B.2 Probabilistic Composition Encoder
where $\mathbf{h}^{0,a}_{i}$ indicates $i$-th row of elemental feature matrix $\mathbf{X}^{a}$, and $l=0,\ldots,L^{\prime}$. After $L^{\prime}$ step message passing steps, we obtain a final representation of composition as follows:
@@
where $\mathbf{W}\in\mathbb{R}^{F\times L^{\prime}F}$ is a learnable weight matrix that reduces the dimension of concatenated representations. Note that $\mathbf{Z}^{a}_{i}$ is equivalent to the $i$-th row of the element embedding matrix $\mathbf{Z}^{a}$ in Equation 2. We also use $L^{\prime}=3$ for composition encoder $f^{a}$. After obtaining the elemental representation matrix $\mathbf{Z}^{a}$, we obtain compositional representation $\mathbf{z}^{a}$ by employing weighted sum pooling, which takes into account the compositional ratio.
Mean $f_{\mu}^{a}$ and Variance $f_{\sigma}^{a}$ Module. After obtaining the elemental representation matrix $\mathbf{Z}^{a}$, we utilize set2set [29] pooling function to obtain the mean vector and diagonal entries of the covariance vector. More specifically, given $\mathbf{Z}^{a}$, we obtain mean vector $\mathbf{z}^{a}_{\mu}$ and diagonal covariance vector $\mathbf{z}^{a}_{\sigma}$ as follows:
$\mathbf{z}^{a}_{\mu}=\mathbf{\hat{z}}^{a}_{\mu}+\mathbf{z}^{a},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{z}}^{a}_{\mu}=\text{Set2set}_{\mu}(\mathbf{Z}^{a}),$ (14)
$\mathbf{z}^{a}_{\sigma}=\mathbf{\hat{z}}^{a}_{\sigma}+\mathbf{z}^{a},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{z}}^{a}_{\sigma}=\text{Set2set}_{\sigma}(\mathbf{Z}^{a}).$ (15)
@@ B.3 Training Details
We also describe the implementation details to enhance the reproducibility. Our method is implemented on Python 3.7.1, PyTorch 1.8.1, and Torch-geometric 1.7.0. All experiments are conducted using a 24GB NVIDIA GeForce RTX 3090. Model hyperparameters are given in Table 7. During training, we clip the gradient to the maximum value of 2 for stability [72].
Table 7: Hyperparameter specifications of PCRL.
@@
Appendix C Additional Experiments
C.1 Representation Space Analysis
Alignment of Composition and Polymorphic Structures. In this part of our study, we delve deeper into the analysis of the representation space for composition and its related polymorphic structures. For a clearer understanding, we’ve utilized t-SNE to visualize the representations of composition that correspond to 8 and 10 polymorphic structures. In the visualizations, each composition is marked by a circle, while the polymorphic structures are denoted by ’x’ markers. In Figure 7, we observe that the composition and its corresponding polymorphic structures are located close to each other, indicating that our composition encoder has effectively learned the information of the polymorphic structures.
Figure 7: Representation space of composition and its corresponding polymorphic structures.
Compositional Representation with Multiple Aspects. In this section, we investigate the various aspects of learned compositional representation. To do so, we first visualize the learned representation of composition in Band Gap dataset using t-SNE [73] in Figure 8. Although PCRL focusing solely on structural information to learn compositional representation, we find that the representation encompasses multiple aspects, including not just the material properties such as band gap (Figure 8 (a)), but also the uncertainties arising from polymorphism (Figure 8 (b)). This implies that PCRL can reliably provide insights into the uncertainties related to polymorphism derived from specific composition, emphasizing its valuable application in material discovery processes.
Figure 8: Visualization of compositional representations in Band Gap dataset.
C.2 Statistical Significance Test
To exhibit the statistical significance of the enhancement, we conducted a paired t-test within the representation learning scenario for each dataset (Table 1). To do so, we additionally conducted 5 separate instances of 5-fold cross-validation, with each instance featuring entirely distinct data splits (amounting to a total of 25 individual runs), and then comparing the means from each set of cross-validation results. The resulting numbers in table represent the p-values from the paired t-tests conducted with PCRL under the following hypotheses: $H_{0}:\mu_{PCRL}=\mu_{Baseline}$, and $H_{1}:\mu_{PCRL}>\mu_{Baseline}$.
Table 8: Statistical significance test results.